Grant is thinking of two numbers. He says that one of the numbers is six times the other number decreased by 2 and the sum of the numbers is 25. What are the numbers?
Answer by mananth(16946)

Grant is thinking of two numbers. He says that one of the numbers is six times the other number decreased by 2 and the sum of the numbers is 25. What are the numbers?
..
let thenumbers be x & y
x+y =25......................1
..
x=6y-2
x-6y=-2......................2
..
multiply equation1 by -1
we get
-x-y=-25
add it to (2)
-6y-y=-25-2
-7y=-27
y=3.85
plug value of y in (1)
x+3.85=25
x=25-3.85
we get x= 21.15
..
m.ananth@hotmail.ca